Organic chemistry help?

2. Eight grams of compound Y was recrystallized us
Answer #1

m = 8 g of Y

V = 103 mL

S = 8/103 = 0.077669 g /mL at T = 100°C

for the second crystallisation:

50 mL --> 0.077669*50 = 3.88345 g of salt should be in solution

then, after crystallisation,

0.957*3.88345 = 3.7164 g were recovered, meaning that

3.88345 -3.7164 = 0.16705 g are in solution

so

S2 = mass/V = 0.16705/50 = 0.003341 g /mL

Solubility at 100°C -->  0.077669 g /mL

Solubility at 20°C --> 0.003341 g /mL